OpenStudy (anonymous):

Use the binomial expression (p + q)n to calculate a binomial distribution with n = 5 and p = 0.3.

OpenStudy (anonymous):

\[ B(n,p) =\binom{n}{k}p^k(1-p)^{n-k} \]The over all sum of any probability distribution is always going to be \(1\). \[ \sum_{k}^nB(n,p) = (p+(1-p))^n=1^n = 1 \]So your question is a bit confusing.

OpenStudy (anonymous):

\[(p+q)^{n}\] thats the actual equations.. but your equations didnt make any sense to me

OpenStudy (anonymous):

By definition \(q=1-p\), so \(p+q=1\).

jimthompson5910 (jim_thompson5910):

I think it wants the individual probabilities that make up the whole distribution, not the sum of all probabilities in the distribution

OpenStudy (anonymous):

Well then how do you find that @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I answered this one a while back, just remembered it http://openstudy.com/users/jim_thompson5910#/updates/516725c4e4b050ab14bf1c31 scroll to the middle/bottom since it's not asked at the top

OpenStudy (anonymous):

@jim_thompson5910 Either it is a probability distribution (which would make sense because it says probability distribution) or it isn't. If it is, then you need something like \(k=3\) or \(k>2\) or some equation to calculate the probability. If it isn't, then the only thing that makes sense if if they give you some \(q\). Either way, not enough information is given.

jimthompson5910 (jim_thompson5910):

yeah i agree it's a bit vague

jimthompson5910 (jim_thompson5910):

but I think they just want the table of the distribution

OpenStudy (anonymous):

OKay thank you guys.. i have one last question and then im done for tonight